Ask
your own question, for FREE!
Mathematics
6 Online
SOME MORE MATH
Still Need Help?
Join the QuestionCove community and study together with friends!
1 o0ver 125
The trick here is it doesn't matter what color she pulls out, even if it's three of the same in a row. There are 5 colors, each with the same chance of being pulled. Any one bead therefore has a 1/5 chance of being pulled. To pull a specific set of beads after returning the ones you pull as Tina does, you would have a much less chance of doing so. To give you an idea, what if you wanted a green bead and then a red one? The green by itself is 1/5, the second by itself is also 1/5. The two together is 1/5*1/5=1/25
So in this case you are doing it three times, which would be 1/5*1/5*1/5
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
3 days ago
3 Replies
2 Medals
6 days ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal